OpenStudy (anonymous):

ABCD-EFGH. The perimeter of ABCD is 32. What is the approximate length of BC?

OpenStudy (anonymous):

A trapezoid problem of similarity EF= 10 FG=12 GH=18 HE=8

OpenStudy (dumbcow):

perimeter of EFGH = 48 ratio = 48/32 = 3/2 assuming BC corresponds to FG 3/2 = 12/BC 3BC = 24 BC = 8